### v1

# On the critical finite-size gap scaling for frustration–free Hamiltonians

Abstract
We prove that the critical finite-size gap scaling for frustration-free Hamiltonians is of inverse-square type. The novelty of this note is that the result is proved on general graphs and for general finite-range interactions. Therefore, the inverse-square critical gap scaling is a robust, universal property of finite-range frustration-free Hamiltonians. This places further limits on their ability to produce conformal field theories in the continuum limit. Our proof refines the divide–and–conquer strategy of Kastoryano and the second author through the refined Detectability Lemma of Gosset-Huang.
1 Setup and main result
1.1 Assumptions on the graph and the Hamiltonian
We consider graphs $\Gamma$ that can be embedded in $\mathbb{R}^{D}$.
Assumption 1.1 (Assumption on the graph).
There exist a dimensional parameter $D\geq 1$, a constant $C_{\Gamma}\geq 1$, and an injective map $\iota:\Gamma\to\mathbb{R}^{D}$ such that
$C_{\Gamma}^{-1}|\iota(i)-\iota(j)|\leq d_{\Gamma}(i,j)\leq C_{\Gamma}|\iota(i)-\iota(j)|,\qquad\textnormal{for all }i,j\in\Gamma.$ (1)
Assumption 1.1 is a rather weak assumption that allows many types of graph structures. The prime example satisfying Assumption 1.1 is the hypercubic lattice $\mathbb{Z}^{D}$ and many standard lattices are also possible (e.g., honeycomb, triangular, fcc, bcc, hcp).
Assumption 1.1 implies that the maximal degree of $\Gamma$ is bounded (see Lemma 3.5 for this fact and other consequences of Assumption 1.1). Examples where Assumption 1.1 is not satisfied include graphs of unbounded degree and the Bethe lattice. While $\Gamma$ is not required to have infinite cardinality, all relevant examples we have in mind have infinite cardinality.
Let $X\subset\Gamma$ be a finite subset. We place at each site of $X$ a $d$-dimensional local Hilbert space. Accordingly, the Hilbert space on the region $X$ is set to
$\mathcal{H}_{X}=\bigotimes_{i\in X}\mathbb{C}^{d}.$
We write $\mathcal{A}_{X}$ for the associated local algebra of bounded linear operators with the natural inclusion $\mathcal{A}_{X}\subset\mathcal{A}_{Y}$ for $X\subset Y$ obtained by taking $A_{X}\otimes\mathrm{Id}_{Y\setminus X}\in\mathcal{A}_{Y}$. To every finite $X$, we associate an interaction $\Phi(X)\in\mathcal{A}_{X}$, a Hermitian bounded linear operator that is subject to the following conditions.
Assumption 1.2 (Assumptions on the interaction).
  • (i)
    $\Phi$ has finite interaction range $R>0$, i.e., $\Phi(X)=0$ whenever $\mathrm{diam}\,(X)>R$.
  • (ii)
    $\Phi(X)\geq 0$ for all $X$ and it is uniformly bounded from above and away from zero, i.e.,
    $\Phi_{\max}:=$ $\sup_{X\subset\Gamma}\|\Phi(X)\|<\infty,$ (2)
    $\Phi_{\min}:=$ $\inf_{\begin{subarray}{c}X\subset\Gamma:\\ \Phi(X)\neq 0\end{subarray}}\operatorname{gap}\Phi(X)>0,$ (3)
    where we introduced $\operatorname{gap}\Phi(X):=\min(\mathrm{spec}\,\Phi(X)\setminus\{0\})$.
Assumption (ii) is a uniformity assumption that allows us to reduce the Hamiltonian to a sum of local projectors. It holds in most of the practically relevant cases, e.g., in any translation-invariant situation. For finite-range interactions, Assumption (ii) is only violated in rather contrived situations, e.g., when the interactions behave very differently at infinity than in the bulk.
Let $\Lambda\subset\Gamma$ be finite. On the Hilbert space $\mathcal{H}_{\Lambda}$, we consider the Hamiltonian
$H_{\Lambda}=\sum_{X\subset\Lambda}\Phi(X)$ (4)
Our main assumption is that the Hamiltonian is frustration-free.
Assumption 1.3 (Assumption on the Hamiltonian).
For every $\Lambda\subset\Gamma$, $H_{\Lambda}$ is frustration-free, i.e., $\ker H_{\Lambda}\neq\{0\}$.
The main quantity of interest is the spectral gap $\operatorname{gap}(\Lambda)$ of $H_{\Lambda}$,
$\operatorname{gap}(H_{\Lambda})=\inf\left(\mathrm{spec}\,H(\Lambda)\setminus\{0\}\right)$ (5)
i.e., its smallest non-zero eigenvalue.
Well-known examples of frustration-free Hamiltonians that can be defined on general graphs $\Gamma$ are the Heisenberg ferromagnet, AKLT Hamiltonians [abdul2020class, affleck1988valence, lucia2023nonvanishing, wei2014hybrid], and Hamiltonians with low-rank interactions [sattath2016local, jauslin2022random, hunter-jones2024gapped]. We recall that almost all mathematical investigations of the size of spectral gaps require frustration-freeness — without this assumption even absolutely fundamental questions like the Haldane conjecture [haldane1983continuum, haldane1983nonlinear] (which asserts that any integer-spin Heisenberg antiferromagnetic chain is gapped) are completely open.
Our contribution in this work is to show that inverse-square critical gap scaling $l^{-2}$ is a robust, universal property of gapless frustration-free Hamiltonians, in the sense that it holds under the rather weak Assumptions 1.1 and 1.2 on the graph and interactions. Our results extend prior ones of this kind [Anshu_2020, gosset2016local, lemm2022quantitatively] which required more restrictive assumptions. It comes with a mild logarithmic correction, which is unimportant for applications.
In particular, our result applies to any boundary condition and excludes gap closing rates that are of inverse length type $l^{-1}$, which would be characteristic of edge modes in conformal field theories [francesco2012conformal, kitaev2006anyons, lemm2019spectral]. Therefore, the result places further limits on the ability of finite-range frustration-free Hamiltonians to produce conformal field theories in the continuum limit.
1.2 Main result
For simplicity, we identify from now on a graph $\Gamma$ satisfying Assumption 1.1 with its Euclidean embedding $\iota(\Gamma)\subset\mathbb{R}^{D}$. This identification can be made without loss of generality, which can be seen by applying $\iota^{-1}$ to the Euclidean rectangles considered below.
We use the following class of domains from [Kastoryano_2018] based on a construction originally appearing in [Cesi_2001]. For each $k\in\mathds{N}$, we let $l_{k}=(3/2)^{k/D}$ and introduce the Euclidean rectangle
$R(k)=[0,l_{k+1}]\times\cdots\times[0,l_{k+D}]\subset\mathbb{R}^{D}.$ (6)
We define $\mathcal{F}_{k}$ to be the collection of subsets of $\Gamma$ which are contained in $R(k)$ up to translations and permutations of the Euclidean coordinates. We denote
$\operatorname{gap}(\mathcal{F}_{k})=\inf_{\Lambda\in\mathcal{F}_{k}}\operatorname{gap}(H_{\Lambda}),\quad\operatorname{gap}(\Gamma)=\inf_{k\geq 1}\operatorname{gap}(\mathcal{F}_{k}).$ (7)
Theorem 1.4 (Main result).
Suppose that
$\operatorname{gap}(\Gamma)=0.$
Then, for every $\epsilon>0$,
$\operatorname{gap}(\mathcal{F}_{k})=o\left(\frac{k^{4+\epsilon}}{l_{k}^{2}}\right),\qquad\textnormal{as }k\to\infty.$ (8)
That is, if the infinite-volume gap vanishes, then the finite-size gaps $\operatorname{gap}(\mathcal{F}_{k})$ “cannot close too slowly”: they must close at least as the inverse square power of the linear dimension, $l_{k}^{-2}$.
We emphasize that the numerator $k^{4+\epsilon}$ is a mild logarithmic correction, since the length parameter $l_{k}=(3/2)^{k/D}$ grows exponentially in $k$.
Similar results to Theorem 1.4 have been proved before using either the divide-and-conquer approach [Kastoryano_2018] or Knabe-type finite-size criteria [gosset2016local, knabe1988energy, lemm2019spectral, lemm2020finite, lemm2022quantitatively]. In particular, an inverse-square gap closing rate was obtained in [Anshu_2020, gosset2016local, lemm2022quantitatively]. These works only apply to nearest-neighbor $2$-body interactions on hypercubic lattices, as well as the honeycomb and triangular lattices. The novelty of the present result is that it covers general graphs and general interactions. Indeed, it covers all finite-range $k$-body interactions. (While additional coarse-graining arguments could be applied to expand the scope of the prior works, they would not give a result of this generality.)
To summarize, the result shows that the inverse-square gap closing speed is a universal property of finite-range frustration-free Hamiltonians. We achieve this generality by using the generality of the divide-and-conquer strategy and amplifying it to near-optimal threshold scaling $\sim l^{-2}$ up to logarithmic corrections via the refined Detectability Lemma due to Gosset-Huang [gosset2016correlation]; see also [anshu2016simple, Anshu_2020, anshu2020entanglement].
Remark 1.5.
The inverse square scaling $l^{-2}$ is well-known to be optimal in view of the Heisenberg ferromagnet, whose spectral gap vanishes as an inverse square on hypercubic lattices in any dimension, which can be easily proved rigorously by constructing explicit spin wave trial states.
2 Proof strategy
2.1 Reduction to sum of projectors
In a preliminary step, we replace each $H_{\Lambda}$ by the sum of projectors
$\tilde{H}_{\Lambda}=\sum_{X\subset\Lambda}h_{X}$ (9)
where $h_{X}$ is the projection onto the range of $\Phi(X)$ (and so $h_{X}=0$ if $\Phi(X)=0$). The fact that $\Phi_{\min}h_{X}\leq\Phi(X)\leq\Phi_{\max}h_{X}$ implies that for any finite $\Lambda\subset\Gamma$,
$\Phi_{\min}\operatorname{gap}(\tilde{H}_{\Lambda})\leq\operatorname{gap}(H_{\Lambda})\leq\Phi_{\max}\operatorname{gap}(\tilde{H}_{\Lambda}).$
Therefore, it suffices to prove Theorem 1.4 for $\tilde{H}_{\Lambda}$. Indeed, the assumption $\operatorname{gap}(H(\Lambda))=0$ implies $\operatorname{gap}(\tilde{H}_{\Lambda})=0$ and the conclusion that $\inf_{\Lambda\in\mathcal{F}_{k}}\operatorname{gap}(\tilde{H}_{\Lambda})=o\left(\frac{k^{4+\epsilon}}{l_{k}^{2}}\right)$ implies $\inf_{\Lambda\in\mathcal{F}_{k}}\operatorname{gap}(H_{\Lambda})=o\left(\frac{k^{4+\epsilon}}{l_{k}^{2}}\right)$. In the following, we always consider the Hamiltonian (9).
2.2 Compatible sequences of domains
Let $\mathcal{F}$ denote a family of finite subsets of $\Gamma$. We recall a result from [Kastoryano_2018], with a minor improvement appearing in [Lucia_2023].
Denote
$\operatorname{gap}(\mathcal{F})=\inf_{\Lambda\in\mathcal{F}}\operatorname{gap}(\Lambda),\qquad\operatorname{gap}(\Lambda)=\operatorname{gap}(\tilde{H}(\Lambda)).$ (10)
Given a finite $A\subset\Gamma$, we write $P_{A}$ for the projection onto the frustration-free ground state of $\tilde{H}_{A}$ (which is identical to the frustration-free ground state of the original Hamiltonian $H_{A}$).
Theorem 2.1 ([Lucia_2023, Theorem 2.3]).
Let $\mathcal{F}$ and $\mathcal{F^{\prime}}$ be two families of finite subsets of a graph $\Gamma$ satisfying Assumption 1.1. Suppose that there exists $s\in\mathds{N}$ and $\delta\in[0,1]$ satisfying the following property: for each $Y\in\mathcal{F}^{\prime}\setminus\mathcal{F}$, there exists $(A_{i},B_{i})_{i=1}^{s}$ pairs of elements in $\mathcal{F}$ such that:
  1. 1.
    $Y=A_{i}\cup B_{i}$ for each $i=1,\dots,s$;
  2. 2.
    $(A_{i}\cap B_{i})\cap(A_{j}\cap B_{j})=\emptyset$ whenever $i\neq j$;
  3. 3.
    $\|P_{A_{i}}P_{B_{i}}-P_{Y}\|\leq\delta$ for every $i=1,\dots,s$.
Then
$\operatorname{gap}(\mathcal{F^{\prime}})\geq\frac{1-\delta}{1+\frac{1}{s}}\operatorname{gap}(\mathcal{F}).$ (11)
The point is that elements of $\mathcal{F}_{k}$ can be nicely decomposed in terms of elements of $\mathcal{F}_{k-1}$.
Let $\alpha\in\{1,\ldots,D\}$. We write $\Pi_{\alpha}$ for the coordinate projection
$\Pi_{\alpha}(x_{1},\ldots,x_{D})=(x_{1},\ldots,x_{\alpha-1},0,x_{\alpha+1},\ldots,x_{D}).$
Proposition 2.2 ([Kastoryano_2018, Proposition 1]).
Let $s_{k}\leq\frac{1}{8}l_{k}$. For each $k\geq 1$, the conditions 1 and 2 of Theorem 2.1 are satisfied for $\mathcal{F}^{\prime}=\mathcal{F}_{k+1}$, $\mathcal{F}=\mathcal{F}_{k}$ and $s=s_{k}$. Moreover, for each $Y\in\mathcal{F}_{k}\setminus\mathcal{F}_{k-1}$, it is possible to choose the pairs $(A_{i},B_{i})_{i=1}^{s_{k}}$ in such a way that
$d_{\Gamma}(A_{i}\setminus B_{i},B_{i}\setminus A_{i})\geq C_{\Gamma}^{-1}\left(\frac{l_{k}}{8s_{k}}-2\right).$ (12)
Moreover, there exists $\alpha\in\{1,\ldots,D\}$ such that
$\Pi_{\alpha}A_{i}=\Pi_{\alpha}B_{i}.$ (13)
Property (13) was not explicitly stated in previous works, but it follows directly from the construction. It is important for us here to connect to the refined Detectability Lemma, which requires us to make a 1D reduction to choose a coarse-graining direction.
Iterating Theorem 2.1 along the sequence $\{\mathcal{F}_{k}\}_{k\geq 1}$ yields the following gap estimate.
Corollary 2.3.
Set
$\delta_{k}:=\sup_{(A,B)\in S_{k}}\|P_{A}P_{B}-P_{A\cup B}\|,$ (14)
where
[…]
Let $k_{0}$ be the smallest $k$ such that $\delta_{k}<1$ for every $k\geq k_{0}$.
Then,
$\operatorname{gap}(\Gamma)\geq\Phi_{\min}\operatorname{gap}(\mathcal{F}_{k_{0}})\prod_{k=k_{0}}^{\infty}\frac{1-\delta_{k}}{1+\frac{1}{s_{k}}}.$ (16)
Note that the constant on the r.h.s. of (16) is strictly positive if and only if $(\delta_{k})_{k}$ and $\left(\frac{1}{s_{k}}\right)_{k}$ are summable sequences. In this case, the “infinite-volume” gap $\inf_{n\geq 1}\operatorname{gap}(\mathcal{F}_{n})$ is lower-bounded by a constant and one commonly says that the infinite system is gapped.
2.3 Proof of Theorem 1.4 assuming Proposition 2.4
The proof of Theorem 1.4 follows from the following Proposition 2.4 whose proof will occupy us in the rest of the paper.
Proposition 2.4.
Let $\lambda_{k}\in(0,1)$ such that
$\operatorname{gap}(\mathcal{F}_{k})\geq\lambda_{k}>0,\quad\forall k.$ (17)
Suppose that there exists a sequence $s_{k}$ such that $(\frac{1}{s_{k}})_{k}$ is summable and
$\liminf_{k\to\infty}\sqrt{\lambda_{k}}\frac{l_{k}}{ks_{k}}>0.$ (18)
Then $\operatorname{gap}(\Gamma)>0$.
Proof of Theorem 1.4.
As explained in Subsection 2.1, it suffices to prove the claim for the Hamiltonians (9). We reason by contradiction. Suppose that
$\operatorname{gap}(\mathcal{F}_{k})\geq c\frac{k^{4+\epsilon}}{l_{k}^{2}},\quad\forall k,$
for some positive constant $c$ and some $\epsilon>0$. Set $\lambda_{k}=c\frac{k^{4+\epsilon}}{l_{k}^{2}}$, which is $<1$ for sufficiently large $k$, and choose $s_{k}=k^{1+\frac{\epsilon}{2}}$, making $(\frac{1}{s_{k}})_{k}$ summable. Moreover
$\liminf_{k\to\infty}\sqrt{\lambda_{k}}\frac{l_{k}}{ks_{k}}=\sqrt{c}\liminf_{k\to\infty}\frac{k^{2+\frac{\epsilon}{2}}}{l_{k}}\frac{l_{k}}{ks_{k}}=\sqrt{c}>0.$
By Proposition 2.4, this implies that $\operatorname{gap}(\Gamma)>0$, leading to a contradiction. ∎
3 Refined overlap bound and proof of Proposition 2.4
We derive an effective upper bound on $\delta_{k}$, defined in (14), in terms of the lower bound on the spectral gap, $\lambda_{k}$. Recall $S_{k}$ defined in (15).
Lemma 3.1 (Refined overlap bound).
There exist $C_{1},C_{2}>0$ such that the following holds. Let $k\geq 1$ and $(A,B)\in S_{k}$. Suppose that $\lambda_{k}\in(0,1)$ satisfies $\operatorname{gap}(\mathcal{F}_{k})\geq\lambda_{k}>0$.
Then,
$\delta_{k}\leq C_{1}\exp\left(-C_{2}\sqrt{\lambda_{k}}\frac{l_{k}}{s_{k}}\right).$ (19)
This is the key improvement over Theorem 11 in [Kastoryano_2018], in which the exponent $\sqrt{\lambda_{k}}$ behaved as $\lambda_{k}$ instead.
We postpone the proof of Lemma 3.1 for the moment. We first combine it with Corollary 2.3 to obtain Proposition 2.4.
Proof of Proposition 2.4 assuming Lemma 3.1.
By Lemma 3.1 and the definition of $\delta_{k}$, we have that
$\delta_{k}\leq\exp\left(-C\sqrt{\lambda_{k}}\frac{l_{k}}{s_{k}}\right).$ (20)
We now apply Corollary 2.3, and in order to show that the r.h.s. of (16) is non-zero, we need to show that $(\delta_{k})_{k}$ is a summable sequence. We will do so by checking that it satisfies the root test:
$\limsup_{k\to\infty}\ \exp\left(-C\sqrt{\lambda_{k}}\frac{l_{k}}{s_{k}}\right)^{1/k}=\exp\left(-C\liminf_{k\to\infty}\sqrt{\lambda_{k}}\frac{l_{k}}{ks_{k}}\right)$
which is strictly smaller than $1$ by assumption. ∎
In the remainder of this section, we prove Lemma 3.1. The proof will follow from a suitable generalization of the refined Detectability Lemma (Lemma 3.6 below), whose setup we describe next. In contrast to [Kastoryano_2018] and rather similarly to [Anshu_2020], we first perform a 1D coarse-graining which is a key ingredient in the refined Detectability Lemma. The differences compared to [Anshu_2020] are that we work in a more general setup (general number of layers and interaction range) and we need to track how Euclidean versus graph geometry enters in the arguments, as both are relevant.
3.1 Coarse-graining
Let $k\geq 1$ and consider $(A,B)\in S_{k}$. We denote $\Lambda=A\cup B$. Let $\alpha\in\{1,\ldots,D\}$ be such that $\Pi_{\alpha}A=\Pi_{\alpha}B$. Without loss of generality, by rotation, we can assume that $\alpha=1$.
We fix a coarse-graining parameter
$t\geq\max\{2,C_{\Gamma}R\}$
and we define the width-$4t$ “columns” centered at $j\in\mathbb{Z}$,
$\mathcal{C}_{j}=([j-2t+1,j+2t-1]\times\mathbb{R}^{D-1})\cap\Lambda,\qquad j\in\mathbb{Z}$
as well as the associated coarse-grained Hamiltonians
$\tilde{H}_{j}=\sum_{X\subset\mathcal{C}_{j}}h_{X}.$
We write $Q_{j}$ for the ground state projector of $\tilde{H}_{j}$.
Definition 3.2.
Let us define two sets of indices:
$\mathcal{I}_{e}(t):=\{(2+6j)\,t\mid j\in\mathds{Z}\},\quad\mathcal{I}_{o}(t):=\{(5+6j)\,t\mid j\in\mathds{Z}\}.$ (21)
Lemma 3.3.
Let #$\#\in\{e,o\}$ and let $m,n\in\mathcal{I}_{\#}(t)$ be distinct. Then $[Q_{m},Q_{n}]=0$.
Proof of Lemma 3.3.
Without loss of generality, let #$\#=e$. Since $m,n\in\mathcal{I}_{e}(t)$, $Q_{m}$ is supported on $[6jt+1,6jt+4t-1]\times\mathbb{R}^{D-1}$ and $Q_{n}$ is supported on $[6j^{\prime}t+1,6j^{\prime}t+4t-1]\times\mathbb{R}^{D-1}$ for two distinct $j,j^{\prime}\in\mathbb{Z}$. The Euclidean distance between these intervals is at least $22t+$ 2. By (1), their $\Gamma$-distance is at least $C_{\Gamma}^{-1}(2t+2)$. Since $t\geq\max\{2,C_{\Gamma}R\}$, we have $C_{\Gamma}^{-1}(2t+2)\geq 2R$. Now $[Q_{m},Q_{n}]=0$ follows because the interaction range is $R$. For #$\#=o$, the argument is analogous. ∎
Lemma 3.3 allows us to group the $Q_{m}$ operators into the even and the odd ones, which commute amongst each other.
Definition 3.4.
The $t$-coarse grained detectability lemma operator $DL(t)$ is given by
$DL(t)=\prod_{m\in\mathcal{I}_{e}(t)}Q_{m}\cdot\prod_{m\in\mathcal{I}_{o}(t)}Q_{m}.$ (22)
3.2 Basic properties of the interaction
We note two basic properties of the interaction $\Phi$ that follow from Assumption 1.2 and that are used in the Detectability Lemma.
Lemma 3.5.
There exist integers $g,L\geq 1$ such that the following holds.
  • $\Phi$ can be decomposed into $L$ layers, i.e., there exists a disjoint decomposition (“$L$-coloring”) of
    $\left\{X\subset\Gamma\;:\;\mathrm{diam}\,(X)\leq R\right\}=\mathcal{X}_{1}\sqcup\mathcal{X}_{2}\sqcup\ldots\sqcup\mathcal{X}_{L}$ (23)
    such that for every $X,Y\in\mathcal{X}_{i}$, it holds that $X\cap Y=\emptyset$.
  • For every subset $X\subset\Gamma$ such that $\Phi(X)\neq 0$, there exist at most $g$ subsets $Y\neq X$ such that $[\Phi(X),\Phi(Y)]\neq 0$.
Proof of Lemma 3.5.
We first observe that Assumption 1.1 implies that, for any $r>0$ and $i\in\Gamma$, the ball
$B^{\Gamma}_{r}(i):=\left\{j\in\Gamma\;:\;d_{\Gamma}(i,j)\leq r\right\}$
has bounded cardinality uniformly in $i$. This follows from the fact that $\iota(B^{\Gamma}_{r}(i))\subset\mathbb{R}^{D}$ is contained in the Euclidean ball $B_{C_{\Gamma}r}(i)$ and all its distinct points are at distance at least $C_{\Gamma}^{-1}>0$.
Concerning the first bullet point, we consider the hypergraph whose hyperedges are those $X\in\mathcal{P}(\Lambda)$ for which $\Phi(X)\neq 0$. We write $\Delta(\Gamma)$ for the degree of the hypergraph, defined as usual as the maximal number of hyperedges intersecting at a vertex. By the finite-range assumption, given any $i\in\Gamma$, the number of $X$ containing $i$ is bounded by $2^{|B_{R}(i)|}$ and therefore this number also bounds the degree. As noted above, the bound on $|B_{R}(i)|$ is uniform in $i$ and so $\Delta(\Gamma)$ is bounded. In this language, $L$ is called the edge-chromatic index of the hypergraph and a simple bound from graph theory is $L\leq\left\lfloor\frac{3\Delta(\Gamma)}{2}\right\rfloor$.
Concerning the second bullet point, we note that $[\Phi(X),\Phi(Y)]\neq 0$ implies $X\cap Y\neq\emptyset$. By the finite-range assumption, $X$ and $Y$ must therefore lie in a $B_{2R}(i)$ for some $i\in\Gamma$. Hence, $g\leq 2^{|B_{2R}(i)|}-1$. As noted above, the bound on $|B_{R}(i)|$ is uniform in $i$ and so $g$ is bounded. ∎
Of course, the bounds on $L$ and $g$ given in the above proof are far from optimal and much better bounds can usually be read off more simply. Since $L$ and $g$ do not enter in the main results, we do not dwell on this issue and refer readers interested in tighter bounds to the combinatorics literature, e.g., [pippenger1989asymptotic] and the recent survey [kang2023graph].
3.3 Refined Detectability Lemma
The following refined Detectability Lemma is due to Gosset and Huang [gosset2016correlation]; see also [anshu2016simple, anshu2020entanglement].
Lemma 3.6 (Refined Detectability Lemma).
Let $k\geq 1$, $(A,B)\in S_{k}$ and $\operatorname{gap}(\mathcal{F}_{k})\geq\lambda_{k}>0$. Denote $\Lambda=A\cup B$. For a normalized state $|\phi\rangle$ orthogonal to the ground state sector of $H_{\Lambda}$, it holds that
$\lVert DL(t)|\phi\rangle\rVert\leq 2\exp\left(-\left(\frac{t}{C_{\Gamma}(L-1)R}-2\right)\sqrt{\frac{\lambda_{k}}{1+g^{2}}}\right)$ (24)
We now prove Lemma 3.6. The first step is to show that one can “smuggle” a polynomial $F$ with $F(0)=1$ and of controlled degree between the products defining $DL(t)$. This polynomial will later be chosen as a suitable Chebyshev polynomial.
Recall the notation (23) for the $L$ layers. The standard detectability lemma operator for $\tilde{H}_{\Lambda}$ from (9) is defined as
$T:=T_{L}\ldots T_{1},\qquad T_{\beta}:=\prod_{X\in\mathcal{X}_{\beta}\cap\Lambda}(1-h_{X}),\quad\beta\in\{1,\ldots,L\}.$ (25)
The following lemma generalizes Lemma 3.1 in [anshu2020entanglement] to any number of layers, interaction range, and Euclidean-embedded graphs.
Lemma 3.7.
Let $F$ be a polynomial of degree at most $\left\lceil\frac{t}{2C_{\Gamma}(L-1)R}-1\right\rceil$, such that $F(0)=1$. Then
$DL(t)=\prod_{m\in\mathcal{I}_{e}(t)}Q_{m}\cdot F\left(\mathds{1}-T^{\dagger}T\right)\cdot\prod_{m\in\mathcal{I}_{o}(t)}Q_{m}.$ (26)
Figure 5 in [anshu2020entanglement] is an instructive depiction of the proof idea in one dimension.
Proof.
For $\beta\in\{1,\ldots,L\}$ and $S\subset\Lambda$, define
$T_{\beta}^{S}=\prod_{\begin{subarray}{c}X\in\mathcal{X}_{\beta}\\ X\cap S\neq\emptyset\end{subarray}}(1-h_{X}).$
Let
$S_{0}=\Lambda\setminus\,\text{supp}\left(\prod_{m\in\mathcal{I}_{e}(t)}Q_{m}\right)$
and define iteratively for $j\geq 1$,
$S_{j}=\left\{i\in\Lambda\;:\;d_{\Gamma}(i,S_{j-1})\leq R\right\}.$
Note that, by construction, $S_{0}\subset S_{1}\subset\cdots\subset S_{n}$ for a suitable $n\geq 1$.
Because of frustration freeness, we have that
$\prod_{m\in\mathcal{I}_{e}(t)}Q_{m}\cdot T_{1}=\prod_{m\in\mathcal{I}_{e}(t)}Q_{m}\cdot T_{1}^{S_{0}}.$
Since each $T_{\beta}$ is a self-adjoint projection, we have that
$T^{\dagger}T=T_{1}T_{2}\ldots T_{L-1}T_{L}T_{L-1}\ldots L_{1},$
and so $(T^{\dagger}T)^{q}=(T_{1}T_{2}\ldots T_{L-1}T_{L}T_{L-1}\ldots T_{2})^{q}T_{1}$. Therefore
$\prod_{m\in\mathcal{I}_{e}(t)}Q_{m}\cdot(T^{\dagger}T)^{q}=\prod_{m\in\mathcal{I}_{e}(t)}Q_{m}\cdot T_{1}^{S_{0}}T_{2}^{S_{1}}\cdots T_{2}^{S_{qL(2L-2)-1}}T_{1}^{S_{q(2L-2)}}.$
Now, if $S_{qL(2L-2)-1}$ is contained in the support of $\prod_{m\in\mathcal{I}_{o}(t)}Q_{m}$, we have that
$T_{1}^{S_{0}}T_{2}^{S_{1}}\cdots T_{2}^{S_{qL(2L-2)-1}}T_{1}^{S_{q(2L-2)}}\cdot\prod_{m\in\mathcal{I}_{o}(t)}Q_{m}=\prod_{m\in\mathcal{I}_{o}(t)}Q_{m}.$
By construction, the set $S_{0}$ is at Euclidean distance $t$ from the complement of the support of $\prod_{m\in\mathcal{I}_{o}(t)}Q_{m}$ and so its graph distance to the complement of the support of $\prod_{m\in\mathcal{I}_{o}(t)}Q_{m}$ is at least $C_{\Gamma}^{-1}t$. By induction, $S_{j}$ is at graph distance $C_{\Gamma}^{-1}t-jR$ from the complement of the support of $\prod_{m\in\mathcal{I}_{o}(t)}Q_{m}$. Therefore, this graph distance is non-zero as long as $j<t/(C_{\Gamma}R)$. Thus by choosing $q\leq\left\lceil t/(2C_{\Gamma}(L-1)R)-1\right\rceil$, we can show that
$\prod_{m\in\mathcal{I}_{e}(t)}Q_{m}\cdot(T^{\dagger}T)^{q}\cdot\prod_{m\in\mathcal{I}_{o}(t)}Q_{m}=DL(t).$
By linearity, we obtain the stated claim. ∎
The identity (26) is used to produce the polynomial and the outside products are then dropped, as summarized in the next corollary. Recall that $P_{\Lambda}$ denotes the projection onto the ground state of $\tilde{H}_{\Lambda}$.
Corollary 3.8.
We have
$\lVert DL(t)P_{\Lambda}^{\perp}\rVert\leq\inf_{F}F_{*}(\lVert TP_{\Lambda}^{\perp}\rVert^{2}),$ (27)
where the infimum runs over polynomials of degree at most $\left\lceil\frac{t}{2C_{\Gamma}(L-1)R}-1\right\rceil$ satisfying $F(0)=1$, and we set
$F_{*}(\epsilon):=\inf_{1-\epsilon\leq x\leq 1}\left|F(x)\right|,\quad\forall\epsilon\in[0,1].$ (28)
Proof of Corollary 3.8.
Let $F$ be any of the admissible polynomials. By Lemma 3.7,
$\lVert DL(t)P_{\Lambda}^{\perp}\rVert\leq\lVert F(\mathds{1}-T^{\dagger}T)P_{\Lambda}^{\perp}\rVert.$
Then
[…]
yields Corollary 3.8. ∎
We recall from [arad2013area] a basic fact about Chebyshev polynomials. We let $T_{q}$ denote the degree-$q$ Chebyshev polynomial.
Lemma 3.9 ([arad2013area]).
For every integer $q\geq 1$ and number $\gamma\in(0,1)$, let
$\mathrm{Step}_{q,\gamma}(x):=\frac{T_{q}\left(\frac{2(1-x)}{1-\gamma}-1\right)}{T_{q}\left(\frac{2}{1-\gamma}-1\right)},\qquad x\in\mathbb{R}.$
Then
$|\mathrm{Step}_{q,\gamma}(x)|\leq 2\exp\left(-2q\sqrt{\gamma}\right),\qquad\gamma\leq x\leq 1.$ (29)
Proof of Lemma 3.6.
Take any vector $|\psi^{\perp}\rangle$ orthogonal to the ground state sector of $H_{\Lambda}$. By the standard Detectability Lemma [anshu2016simple, Corollary 1], we have
$\|T|\psi^{\perp}\rangle\|^{2}\leq\frac{1}{1+\lambda_{k}/g^{2}}$
and therefore $\|TP_{\Lambda}^{\perp}\|^{2}\leq\frac{1}{1+\lambda_{k}/g^{2}}$.
By applying Corollary 3.8 and Lemma 3.9 with $q=\left\lceil\frac{t}{2C_{\Gamma}(L-1)R}-1\right\rceil$ and $\gamma=\frac{\lambda_{k}}{\lambda_{k}+g^{2}}\in(0,1)$, we obtain
$\|DL(t)|\phi\rangle\|\leq\inf_{F}F_{*}(\|LP_{\Lambda}^{\perp}\|^{2})\leq$ $2\exp\left(-2\left\lceil\frac{t}{2C_{\Gamma}(L-1)R}-1\right\rceil\sqrt{\frac{\lambda_{k}}{\lambda_{k}+g^{2}}}\right)$
$\leq$ $2\exp\left(-\left(\frac{t}{C_{\Gamma}(L-1)R}-2\right)\sqrt{\frac{\lambda_{k}}{1+g^{2}}}\right)$
where we used $\lambda_{k}\leq 1$ in the last step. This proves Lemma 3.6. ∎
3.4 Proof of Lemma 3.1
Let $k\geq 1$, $(A,B)\in S_{k}$, and denote $\Lambda=A\cup B$. We perform the coarse-graining as described in Subsection 3.1.
We now define $M_{A}$ and $M_{B}$ by reordering the projector $Q_{m}$ appearing in $DL(t)$ in such a way that $DL(t)=M_{A}M_{B}$. We set
$M_{A}=\prod_{\begin{subarray}{c}m\in\mathcal{I}_{e}(t):\\ \mathrm{supp}\,Q_{m}\cap A\setminus B\neq\emptyset\end{subarray}}Q_{m}\prod_{\begin{subarray}{c}m\in\mathcal{I}_{o}(t):\\ \exists m^{\prime}\in\mathcal{I}_{e}(t):\\ \mathrm{supp}\,Q_{m^{\prime}}\cap A\setminus B\neq\emptyset,\\ \mathrm{supp}\,Q_{m}\cap\mathrm{supp}\,Q_{m^{\prime}}\neq\emptyset\end{subarray}}Q_{m}$
and we write $M_{B}$ for the remaining $Q_{m}$ appearing in $DL(t)$, ordered in the same way as they appear in $DL(t)$. Then $DL(t)=M_{A}M_{B}$ because, in $M_{A}$, the odd-type terms in the second product commute precisely with all even-type terms except for those in the first product. If the Euclidean distance between $A\setminus B$ and $B\setminus A$ exceeds $8t$, then $M_{A}$ is supported on $A$ and $M_{B}$ is supported on $B$.
Recall that $P_{A}$ denotes the projection onto the ground state of $\tilde{H}_{A}$. We have that $P_{A}=P_{A}M_{A}$, and therefore $P_{A}M_{B}=P_{A}M_{A}M_{B}=P_{A}DL(t)$. This implies that
$\lVert(P_{A}-M_{A})M_{B}\rVert=\lVert DL(t)P_{A}^{\perp}\rVert\leq\lVert DL(t)P_{AB}^{\perp}\rVert,$
where the inequality follows from the fact that $P_{A}^{\perp}\leq P_{AB}^{\perp}$, a consequence of frustration-freeness. Since it also holds that $P_{B}=P_{B}M_{B}$ and $P_{B}M_{A}=P_{B}DL(t)$, the analogous bound with $A$ and $B$ interchanged. This gives that
$\lVert P_{A}P_{B}-P_{AB}\rVert=$ $\lVert P_{A}P_{B}-M_{A}P_{B}\rVert+\lVert M_{A}P_{B}-M_{A}M_{B}\rVert+\lVert M_{A}M_{B}-P_{AB}\rVert$
$\leq$ $\lVert(P_{A}-M_{A})P_{B}\rVert+\lVert M_{A}(P_{B}-M_{B})\rVert+\lVert DL(t)P_{AB}^{\perp}\rVert\leq 3\lVert DL(t)P_{AB}^{\perp}\rVert.$
By Lemma 3.6, we have
$\lVert DL(t)P_{AB}^{\perp}\rVert\leq 2\exp\left(-\left(\frac{t}{C_{\Gamma}(L-1)R}-2\right)\sqrt{\frac{\lambda_{k}}{1+g^{2}}}\right).$
By (12) and Assumption 1.1, we can choose $t$ as a constant multiple of $\frac{l_{k}}{s_{k}}$. Defining $C_{1},C_{2}>0$ appropriately, this proves Lemma 3.1. ∎
Acknowledgments
The authors thank the organizers of the online Lattice Seminar, during which the idea for this work arose. The research of M.L. is supported by the DFG through the grant TRR 352 – Project-ID 470903074. A. L. acknowledges financial support from grants PID2020-113523GB-I00 and CEX2023-001347-S, funded by MICIU/AEI/ 10.13039/501100011033, and from grant RYC2019-026475-I funded by MICIU/AEI/ 10.13039/501100011033 and “ESF Investing in your future”.
[…]